OpenStudy (anonymous):

What is the probability of picking a red card greater than 9 from a standard deck of 52 playing cards? (Assume that aces =1.)

OpenStudy (anonymous):

read card grater than 9 : 10 J Q K A 5 x2(red card)=10 probabiliy=10/52=5/26

OpenStudy (danjs):

10 , J, Q, K 2 each (diamond, heart)

OpenStudy (danjs):

Aces = 1 , and do not count here as being greater than 9